# Simpler Diffusion (SiD2): 1.5 FID on ImageNet512 with pixel-space diffusion

### v1

@@ Abstract
Latent diffusion models have become the popular choice for scaling up diffusion models for high resolution image synthesis. Compared to pixel-space models that are trained end-to-end, latent models are perceived to be more efficient and to produce higher image quality at high resolution. Here we challenge these notions, and show that pixel-space models can in fact be very competitive to latent approaches both in quality and efficiency, achieving 1.5 FID on ImageNet512 and new SOTA results on ImageNet128 and ImageNet256.
We present a simple recipe for scaling end-to-end pixel-space diffusion models to high resolutions. 1: Use the sigmoid loss (Kingma & Gao, 2023) with our prescribed hyper-parameters. 2: Use our simplified memory-efficient architecture with fewer skip-connections. 3: Scale the model to favor processing the image at high resolution with fewer parameters, rather than using more parameters but at a lower resolution. When combining these three steps with recently proposed tricks like guidance intervals, we obtain a family of pixel-space diffusion models we call Simple Diffusion v2 (SiD2).
@@ 1 Introduction
Figure 1: End-to-end diffusion performance over the years, on ImageNet512.
Diffusion models have shown to be tremendously successful in domains such as image, video and audio generation. An important technique has been latent diffusion (Rombach et al., 2022), which has allowed scaling to large resolution images where traditional diffusion models have failed.
In EDM2 (Karras et al., 2023), the performance characteristics of latent diffusion model on ImageNet512 are shown, which is a benchmark for which reliable metrics exist. They show that compared to other approaches, latent diffusion in combination with their architecture choices have superior scaling behaviour, leading to an impressive FID of 1.40 on ImageNet512 with guidance intervals. The performance differences are even more drastic when compared to pixel-space diffusion approaches. Why is pixel-space diffusion so difficult to train on high resolution images?
Latent models have their own challenges: It can be difficult to train the autoencoders which require regularization hyperparameters (Rombach et al., 2022). Further, the score of multiple diffusion models can be used in an ensemble, but this only works if they are trained using the same autoencoder. Finally, underfitting may result in out of distribution latents, which will cause the decoder to generate artifacts. A single-stage pixel diffusion model does not have these issues.
Over the years, end-to-end pixel-space diffusion has improved considerably (Figure 1). The main problem was that the schedule and loss of diffusion models work well for small resolution images, for example $32^{2}$ or $64^{2}$. However, for larger resolutions the schedules and loss can be misspecified (Chen, 2023; Hoogeboom et al., 2023). It was further shown (Kingma & Gao, 2023) that it can be helpful to disentangle loss weighting and time sampling for more loss control and even better performance. However, the best ImageNet512 known to date for pixel-space diffusion is still only 2.65.
The main contribution of our paper is to improve pixel-space diffusion on high resolution images while keeping as close to the original formulation of diffusion models as possible. Specifically, we analyze and tune the sigmoid loss weighting from (Kingma & Gao, 2023), and show why it should be preferred over their favored EDM monotonic. We further propose architectural improvements that lead to even better performance, namely flop heavy scaling and removing block skip-connections (Residual U-ViTs). Most notably, we improve the state-of-the-art of pixel-space diffusion from 2.65 to 1.50 on ImageNet512 (see also Figure 1). Further, we also achieve overall state-of-the-art on ImageNet128 and ImageNet256 without heavy tuning.
@@ 1.1 Background
which defines a distribution $q({\bm{z}}_{t}|{\bm{x}})$ such that ${\bm{z}}_{0}$ is close to clean data and ${\bm{z}}_{1}$ is close to Gaussian noise. In this paper we assume a variance-preserving (Song et al., 2021) schedule, which is defined as $\alpha_{t}=\sqrt{\sigma(\lambda_{t})}$ and $\sigma_{t}=\sqrt{\sigma(-\lambda_{t})}$, with $\lambda_{t}=\log\big{(}\alpha_{t}^{2}/\sigma_{t}^{2}\big{)}$ being the log signal-to-noise ratio, which monotonically decreases with time as $t\rightarrow 1$, and $\sigma(\cdot)$ being the sigmoid function. It can be shown that other types of schedules such as the variance-exploding schedule or the flow matching schedule (Lipman et al., 2023) can always be converted to a variance-preserving schedule with a linear scaling of ${\bm{z}}_{t}$. The key defining property of a noise schedule is the log signal-to-noise ratio $\lambda_{t}$. A neural network is used to predict $\hat{{\bm{x}}}=\hat{{\bm{x}}}_{\theta}({\bm{z}}_{t},t)$ that effectively estimates $\mathbb{E}[{\bm{x}}|{\bm{z}}_{t}]$, which can be done by minimizing a weighted mean squared error loss:
@@
where $w(\lambda_{t})$ is the weighting function. ${\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}-\mathrm{d}\lambda_{t}/\mathrm{d}t}$ cancels out the effect of the noise schedule $\lambda_{t}$ on the loss. In other words, $L({\mathbf{x}})$ can be rewritten as $\int-w(\lambda)\|{\bm{x}}-\hat{{\bm{x}}}\|^{2}d\lambda$ that is invariant to the choice of $\lambda_{t}$. However, $\lambda_{t}$ still impacts the sampling efficiency of the Monte Carlo estimator of the expected loss, and thus affects the training dynamics of diffusion models. This decouples noise schedule $\lambda_{t}$ and weighting function $w(\lambda)$ as a function of $\lambda$ (Kingma & Gao, 2023).
The above squared error loss can be derived from the variational interpretation of diffusion models (Kingma et al., 2021). Given two time steps $s<t$, we define the denoising distribution $p({\bm{z}}_{s}|{\bm{z}}_{t})$ in the same formulation as the true posterior $q({\bm{z}}_{s}|{\bm{z}}_{t},{\bm{x}})$, but replace the ground truth ${\bm{x}}$ with the learned approximation $\hat{{\bm{x}}}\approx\mathbb{E}[{\bm{x}}|{\bm{z}}_{t}]$:
$p({\bm{z}}_{s}|{\bm{z}}_{t})=q({\bm{z}}_{s}|{\bm{z}}_{t},{\bm{x}}=\hat{{\bm{x}}})=\mathcal{N}({\bm{z}}_{s}|\mu_{t\to s},\sigma_{t\to s}),$ (3)
where ${\sigma_{t\to s}^{2}=\Big{(}1/\sigma_{s}^{2}+\alpha_{t|s}^{2}/\sigma_{t|s}^{2}\Big{)}^{-1}}$ and ${\mu_{t\to s}=\sigma_{t\to s}^{2}\Big{(}\alpha_{t|s}/\sigma_{t|s}^{2}{\bm{z}}_{t}+\alpha_{s}/\sigma_{s}^{2}{\bm{x}}\Big{)}}$ with $\alpha_{t|s}=\alpha_{t}/\alpha_{s}$ and $\sigma_{t|s}^{2}=\sigma_{t}^{2}-\alpha_{t|s}\sigma_{s}^{2}$. Note that in practice during sampling, the standard deviation for $p({\bm{z}}_{s}|{\bm{z}}_{t})$ is chosen to be a log-linear interpolation between $\sigma_{t|s}$ and $\sigma_{t\to s}$ of the form $\sigma_{t\to s}^{\gamma}\sigma_{t|s}^{1-\gamma}$ where $\gamma=0.3$ was set empirically, which is important as FID is sensitive to over-smooth or grainy images.
@@ 2.1 Revisiting the sigmoid loss
Recall that one of the proposed losses in VDM++ (Kingma & Gao, 2023) is the sigmoid loss. They expressed the loss with respect to ${\bm{{\epsilon}}}$-mse: $L({\bm{x}})=\mathbb{E}_{t\sim\mathcal{U}(0,1)}{-\frac{\mathrm{d}\lambda_{t}}{\mathrm{d}t}}\tilde{w}(\lambda_{t})\|{\bm{{\epsilon}}}-\hat{{\bm{{\epsilon}}}}\|^{2}$, which is connected to ${\bm{x}}$-mse through $\|{\bm{{\epsilon}}}-\hat{{\bm{{\epsilon}}}}\|^{2}=\exp(\lambda_{t})\|{\bm{x}}-\hat{{\bm{x}}}\|^{2}$ and $\tilde{w}(\lambda_{t})=\exp(\lambda_{t})w(\lambda_{t})$. In VDM++, viewing the loss in ${\bm{{\epsilon}}}$-mse enables a natural connection between the weighted loss and the ELBO and showed that losses should be monotonically decreasing on ${\bm{{\epsilon}}}$-mse. Hence, they propose the sigmoid loss: $\tilde{w}(\lambda_{t})=\sigma(b-\lambda_{t})$, where $b$ denotes the bias or shift and $\sigma$ is the sigmoid function. However, they ultimately opted for another loss named EDM-monotonic for high resolution image generation. See Figure 3(a) for a comparison.
@@
$\sigma(b-\lambda_{t})\|{\bm{{\epsilon}}}-\hat{{\bm{{\epsilon}}}}\|^{2}=\sigma(b-\lambda_{t})\exp(\lambda_{t})\|{\bm{x}}-\hat{{\bm{x}}}\|^{2}={\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\exp(-b)}\sigma(\lambda_{t}-b)\|{\bm{x}}-\hat{{\bm{x}}}\|^{2}.$ (4)
@@
Studying both the form of the sigmoidal and EDM-monotonic weighting functions in Figure 3, we observe that both weighting functions are monotonically decreasing in ${\bm{{\epsilon}}}$-mse, but only the sigmoidal weighting function is monotonically increasing in ${\bm{x}}$-mse. Therefore, we expect the sigmoidal weighting to be a better candidate to scale to high resolution generation than the EDM-monotonic version.
@@ 2.2 Flop heavy scaling
When one deals with a limited data budget, it can be difficult to balance scaling of models and regularization. Our small model version at $512\times 512$ resolution starts out with 4 $\times$ 4 patching, and subsequently runs a U-ViT (Hoogeboom et al., 2023) on the 128 $\times$ 128 $\times$ 48 input (assuming the raw input is of $3$ channels). To make the model more expressive, one of the most common options is to double the number of channels for every layer, leading to 2×2$\times$2 × model size. However, our experiments show that this results in diminishing returns. We hypothesize that it is because the model suffers from over-fitting when trained with limited data budget such as ImageNet. Indeed, when we train using distribution augmentation (Jun et al., 2020), the best performance improves from 1.9 to 1.6. This however suggests another way to scale the model: Keep the architecture and parameter count almost the same and instead run on 2 $\times$ 2 patched input (so 256 $\times$ 256 $\times$ 12 for 512 inputs). Similar to 2×2$\times$2 × model size, reducing the patching size also increases FLOPS by roughly a factor of 4. However, we found this model performs even better to about 1.5 without the need of distribution augmentation. We call this model flop heavy, because it has a much higher flop over parameter ratio.
Flop heavy scaling has another interesting property. Suppose we have a small model trained to fit input of size $256\times 256$ with $2\times 2$ patching. Using the flop heavy scaling at $512\times 512$ resolution enables us to directly finetune from the small model trained at $256\times 256$ with exactly the same amount of parameters, whereas its counterpart version at $512\times 512$ resolution and with $4\times 4$ patching needs to introduce additional parameters to handle the more input channels. This method of finetuning to higher resolutions has already been popularized in the context of text-to-image generation (Rombach et al., 2022). Here we find that flop heavy scaling is a more favorable way to scale on smaller datasets such as ImageNet. Of course, in principle one could also train a model with more channels and $4\times 4$ patching on a lower resolution, but this is not done in practice to train ‘small’ models.
@@ 2.3 Residual U-ViTs: Removing blockwise skip-connections
Figure 4: Residual U-ViT. Instead of blockwise skip-connections, every level has a single skip-connection. If the neural network blocks are initialized to be identity, the entire network is also identity at initialization.
U-Nets require that sub-sampled levels (for example via average pooling) are bypassed by skip-connections. However, this makes the architecture very different from large-scale language models such as transformers, where these skip-connections are not used and residual connections suffice. In this paper we use the U-ViT model of Hoogeboom et al. (2023), which is a variant of U-Nets with self-attention layers (with convolutional layers) being replaced by transformer blocks. Hence, that part of the architecture is closer to standard transformers.
In this paper, we further simplify model definition with Residual U-ViTs. The blockwise skip-connections are removed, and in their place only a single skip-connection is used per downsampling operation. To be precise, let $f_{d},f_{u}$ denote a down and up level of a U-ViT and let $f_{m}$ denote the middle stage. This two-stage Residual U-ViT can now be described by:
@@
where $\operatorname{D}$ is typically an average pooling and linear and $\operatorname{U}$ is typically a linear and nearest neighbour upsampling. Defining this network to preserve an identity function at initialization is now much easier: As long as $f_{d}$, $f_{u}$ and $f_{m}$ are initialized to be identity, the entire $f$ is also identity at initialization. The generalization to any-depth U-ViTs is also straightforward. For example, we can define another two levels of U-ViTs in place of $f_{m}$ recursively, where the nested U-ViTs can also be guaranteed to be initialized as an identity function.

### v2

@@ Abstract
Latent diffusion models have become the popular choice for scaling up diffusion models for high resolution image synthesis. Compared to pixel-space models that are trained end-to-end, latent models are perceived to be more efficient and to produce higher image quality at high resolution. Here we challenge these notions, and show that pixel-space models can be very competitive to latent models both in quality and efficiency, achieving 1.5 FID on ImageNet512 and new SOTA results on ImageNet128, ImageNet256 and Kinetics600.
We present a simple recipe for scaling end-to-end pixel-space diffusion models to high resolutions. 1: Use the sigmoid loss-weighting [25] with our prescribed hyper-parameters. 2: Use our simplified memory-efficient architecture with fewer skip-connections. 3: Scale the model to favor processing the image at a high resolution with fewer parameters, rather than using more parameters at a lower resolution. Combining these with guidance intervals, we obtain a family of pixel-space diffusion models we call
Simpler Diffusion (SiD2).
@@ 1 Introduction
Diffusion models have been tremendously successful in domains such as image, video and audio generation. For high resolution data, latent diffusion models [32] have been particularly successful: Through encoding the data in a lower dimensional space, these models sidestep the difficulties faced by more traditional pixel-space diffusion models in scaling up to high resolution. This makes latent diffusion models also scale better in training compute than other approaches, as shown by [20].
@@
Figure 1: End-to-end pixel diffusion performance on ImageNet512.
@@
However, latent models have their own challenges: It can be difficult to train the autoencoders that produce the latents, which in practice requires careful tuning of regularization hyperparameters [32]. Even when trained well, the decoder of a latent diffusion model can produce artifacts or otherwise imperfect reconstructions, especially when it is applied to latents that are out of distribution compared to the training set. A single-stage pixel-space diffusion model does not have these issues, and so we ask the question: Can pixel-space diffusion models scale as well as latent models with the right architecture and hyperparameters?
Over the years, end-to-end pixel-space diffusion has improved considerably (Figure 1). The main problem was that the standard schedule and loss-weighting [11] work well for small resolution images, like $32^{2}$ or $64^{2}$, but not at higher resolution [3, 14]. It was further shown [25] that it can be helpful to disentangle loss weighting and time sampling for improved control over both, leading to better performance at higher resolution. However, in previous works pixels-space models still did not match the best latent models on standard benchmarks like ImageNet512.
The main contribution of this paper is to improve pixel-space diffusion on high resolution images while keeping as close to the original formulation of diffusion models as possible. Specifically, we analyze and tune the sigmoid loss weighting from [25], and show why it should be preferred over their favored EDM monotonic. We further propose architectural improvements that lead to even better performance, namely flop heavy scaling and removing block skip-connections (Residual U-ViTs). Most notably, we improve the state-of-the-art of pixel-space diffusion from 2.65 to 1.48 on ImageNet512 (see also Figure 1). Further, we also achieve overall state-of-the-art on ImageNet128, ImageNet256 and Kinetics600 without heavy tuning.
@@ 1.1 Background
which defines a distribution $q({\bm{z}}_{t}|{\bm{x}})$ such that ${\bm{z}}_{0}$ is close to clean data and ${\bm{z}}_{1}$ is close to Gaussian noise. In this paper we assume a variance-preserving [41] schedule, which is defined as $\alpha_{t}=\sqrt{\sigma(\lambda_{t})}$ and $\sigma_{t}=\sqrt{\sigma(-\lambda_{t})}$, with $\lambda_{t}=\log\big{(}\alpha_{t}^{2}/\sigma_{t}^{2}\big{)}$ being the log signal-to-noise ratio, which monotonically decreases with time as $t\rightarrow 1$, and $\sigma(\cdot)$ being the sigmoid function. It can be shown that other types of schedules such as the variance-exploding schedule or the flow matching schedule [28] can always be converted to a variance-preserving schedule with a linear scaling of ${\bm{z}}_{t}$ [25]. The key defining property of a noise schedule is the log signal-to-noise ratio $\lambda_{t}$. A neural network is used to predict $\hat{{\bm{x}}}=\hat{{\bm{x}}}_{\theta}({\bm{z}}_{t},t)$ that effectively estimates $\mathbb{E}[{\bm{x}}|{\bm{z}}_{t}]$, which can be done by minimizing a weighted mean squared error loss:
@@
where $w(\lambda_{t})$ is the weighting function. ${\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}-\mathrm{d}\lambda_{t}/\mathrm{d}t}$ cancels out the effect of the noise schedule $\lambda_{t}$ on the loss. In other words, $L({\mathbf{x}})$ can be rewritten as $\int-w(\lambda)\|{\bm{x}}-\hat{{\bm{x}}}\|^{2}d\lambda$ that is invariant to the choice of $\lambda_{t}$. However, $\lambda_{t}$ still impacts the sampling efficiency of the Monte Carlo estimator of the expected loss, and thus affects the training dynamics of diffusion models. This decouples noise schedule $\lambda_{t}$ and weighting function $w(\lambda)$ as a function of $\lambda$ [25].
The above squared error loss can be derived from the variational interpretation of diffusion models [26]. Given two time steps $s<t$, we define the denoising distribution $p({\bm{z}}_{s}|{\bm{z}}_{t})$ in the same formulation as the true posterior $q({\bm{z}}_{s}|{\bm{z}}_{t},{\bm{x}})$, but replace the ground truth ${\bm{x}}$ with the learned approximation $\hat{{\bm{x}}}\approx\mathbb{E}[{\bm{x}}|{\bm{z}}_{t}]$:
$\small p({\bm{z}}_{s}|{\bm{z}}_{t})=q({\bm{z}}_{s}|{\bm{z}}_{t},{\bm{x}}=\hat{{\bm{x}}})=\mathcal{N}({\bm{z}}_{s}|\mu_{t\to s},\sigma_{t\to s}),$ (3)
where $\small{\sigma_{t\to s}^{2}=\Big{(}1/\sigma_{s}^{2}+\alpha_{t|s}^{2}/\sigma_{t|s}^{2}\Big{)}^{-1}}$ and $\small{\mu_{t\to s}=\sigma_{t\to s}^{2}\Big{(}\alpha_{t|s}/\sigma_{t|s}^{2}{\bm{z}}_{t}+\alpha_{s}/\sigma_{s}^{2}{\bm{x}}\Big{)}}$ with $\alpha_{t|s}=\alpha_{t}/\alpha_{s}$ and $\sigma_{t|s}^{2}=\sigma_{t}^{2}-\alpha_{t|s}\sigma_{s}^{2}$. Note that during sampling, the standard deviation for $p({\bm{z}}_{s}|{\bm{z}}_{t})$ is chosen to be a log-linear interpolation between $\sigma_{t|s}$ and $\sigma_{t\to s}$ of the form $\sigma_{t\to s}^{\gamma}\sigma_{t|s}^{1-\gamma}$ where $\gamma=0.3$ was set empirically, which is important as FID is sensitive to over-smooth or grainy images.
@@ 2.1 Revisiting the sigmoid loss
Recall that one of the proposed losses in VDM++ [25] is the sigmoid loss. They expressed the loss with respect to ${\bm{{\epsilon}}}$-mse: $L({\bm{x}})=\mathbb{E}_{t\sim\mathcal{U}(0,1)}{-\frac{\mathrm{d}\lambda_{t}}{\mathrm{d}t}}\tilde{w}(\lambda_{t})\|{\bm{{\epsilon}}}-\hat{{\bm{{\epsilon}}}}\|^{2}$, which is connected to ${\bm{x}}$-mse through $\|{\bm{{\epsilon}}}-\hat{{\bm{{\epsilon}}}}\|^{2}=\exp(\lambda_{t})\|{\bm{x}}-\hat{{\bm{x}}}\|^{2}$ and $\tilde{w}(\lambda_{t})=\exp(\lambda_{t})w(\lambda_{t})$. In VDM++, viewing the loss in ${\bm{{\epsilon}}}$-mse enables a natural connection between the weighted loss and the ELBO and showed that losses should be monotonically decreasing on ${\bm{{\epsilon}}}$-mse. Hence, they propose the sigmoid loss: $\tilde{w}(\lambda_{t})=\sigma(b-\lambda_{t})$, where $b$ denotes the bias or shift and $\sigma$ is the sigmoid function. However, they ultimately opted for another loss-weighting named EDM-monotonic for high resolution image generation, see Figure 3(a).
@@
$\begin{split}\sigma(b-\lambda_{t})\|{\bm{{\epsilon}}}-\hat{{\bm{{\epsilon}}}}\|^{2}&=\sigma(b-\lambda_{t})\exp(\lambda_{t})\|{\bm{x}}-\hat{{\bm{x}}}\|^{2}\\ &={\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\exp(b)}\sigma(\lambda_{t}-b)\|{\bm{x}}-\hat{{\bm{x}}}\|^{2}.\end{split}$ (4)
@@
Studying both the form of the sigmoid and EDM-monotonic weighting functions in Figure 3, we observe that both weighting functions are monotonically decreasing in ${\bm{{\epsilon}}}$-mse, but only the sigmoid weighting is monotonically increasing in ${\bm{x}}$-mse. Therefore, we expect the sigmoid weighting to be a better candidate to scale to high resolution generation than the EDM-monotonic version.
@@ 2.2 Flop heavy scaling
When one deals with a limited data budget, it can be difficult to balance scaling of models and regularization. Our small model version at $512\times 512$ resolution starts out with 4 $\times$ 4 patching, and subsequently runs a U-ViT [14] on the 128 $\times$ 128 $\times$ 48 input (assuming the raw input is of $3$ channels). To make the model more expressive, one of the most common options is to double the number of channels for every layer, leading to 2×2$\times$2 × model size. However, our experiments show that this results in diminishing returns. We hypothesize that it is because the model suffers from over-fitting when trained with limited data budget such as ImageNet. Indeed, when we train using distribution augmentation [18], the best performance improves from 1.9 to 1.6. This however suggests another way to scale the model: Keep the architecture and parameter count almost the same and instead run on 2 $\times$ 2 patched input (so 256 $\times$ 256 $\times$ 12 for 512 inputs). Similar to 2×2$\times$2 × model size, reducing the patching size also increases FLOPS by roughly a factor of 4. However, we found this model performs even better to about 1.5 without the need of distribution augmentation. We call this model flop heavy, because it has a much higher flop over parameter ratio.
Flop heavy scaling has another interesting property. Suppose we have a small model trained to fit input of size $256\times 256$ with $2\times 2$ patching. Using the flop heavy scaling at $512\times 512$ resolution enables us to directly finetune from the small model trained at $256\times 256$ with exactly the same amount of parameters, whereas its counterpart version at $512\times 512$ resolution and with $4\times 4$ patching needs to introduce additional parameters to handle the more input channels. This method of finetuning to higher resolutions has already been popularized in the context of text-to-image generation [32]. Here we find that flop heavy scaling is a more favorable way to scale on smaller datasets such as ImageNet. Of course, in principle one could also train a model with more channels and $4\times 4$ patching on a lower resolution, but this is not done in practice to train ‘small’ models.
@@ 2.3 Residual U-ViTs: Removing blockwise skip-connections
U-Nets require that sub-sampled levels (for example via average pooling) are bypassed by skip-connections. However, this makes the architecture very different from large-scale language models such as transformers, where these skip-connections are not used and residual connections suffice.
In this paper we use the U-ViT model of [14], which is a variant of U-Nets with self-attention layers (with convolutional layers) being replaced by transformer blocks. Hence, that part of the architecture is closer to standard transformers.
In this paper, we further simplify model definition with Residual U-ViTs, see Figure 4. The blockwise skip-connections are removed, and in their place only a single skip-connection is used per downsampling operation. To be precise, let $f_{d},f_{u}$ denote a down and up level of a U-ViT and let $f_{m}$ denote the middle stage. This two-stage Residual U-ViT can now be described by:
@@
where $\operatorname{D}$ is typically an average pooling and linear and $\operatorname{U}$ is typically a linear and nearest neighbour upsampling. Defining this network to preserve an identity function at initialization is now much easier: As long as $f_{d}$, $f_{u}$ and $f_{m}$ are initialized to be identity, the entire $f$ is also identity at initialization. The generalization to any-depth U-ViTs is also straightforward:we can define another two levels of U-ViTs in place of $f_{m}$ recursively, where the nested U-ViTs can also be guaranteed to be initialized as an identity function.
@@
Figure 4: Residual U-ViT with levelwise skip-connections (instead of blockwise). When the network blocks are initialized to be identity, the entire network is also an identity function at initialization.